the perimeter of a standard sized rectangular rug is 44 feet. the width is 2 feet less than the length . find the dimension
Answer by ewatrrr(24785)


Hi,
Let x represent the length, then the width is (x-2)
Question states
2*x + 2(x-2) = 44ft (the perimeter)
solving for x
4x -4 = 44ft
4x = 48ft
x = 12ft, the length, width is 10ft (12-2)
checking our answer
2*12 + 2*10 = 44